40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue May 19 19:24:30 2015, rana, Update, PEM, Seismic BLRMS filters BLRMS_BP.pdfBLRMS_imp.pdf
    Reply  Mon Sep 7 19:46:46 2015, rana, Update, PEM, Seismic BLRMS filters T240_150907.png
       Reply  Wed Sep 9 15:45:11 2015, ericq, Update, PEM, Inverted STS filters in C1OAF 
Message ID: 11582     Entry time: Mon Sep 7 19:46:46 2015     In reply to: 11316     Reply to this: 11587
Author: rana 
Type: Update 
Category: PEM 
Subject: Seismic BLRMS filters 

As it turned out, the "STS" BLRMS filters were all a mess, so I fixed them up today:

  • BP and LP filters were non-existent for the 2 low frequency bands: 0.01-0.03 & 0.03-0.1 Hz. The 0.01-0.03 is just seeing tilt noise (its big in X & Y, but not in Z), but the other band is able to cleanly see the primary microseism at 0.06 Hz.
  • There was some mixup and some BP filter banks had low pass filters while one of the LP banks had a BP filter.
  • There were different filters between the X, Y & Z directions.
  • The low pass filters had enough ringing in the impulse response that their outputs could sometimes go negative and make the SQRT block output NaN.

After tuning:

  • All bandpass filters are 4th order Butterworth bandpass with the corners at the band edges (e.g. 1- 3  Hz)
  • All low pass are the same, just scaled by the frequency band. They have a pair of real poles and a pair of 35 deg poles. The pole frequencies are set so that there is 40 dB of attenuation at twice the frequency of the low end of the bandpass. i.e. for the 1-3 Hz band, the low pass has > 40 dB atten at 2 Hz.
  • The 3-10 and 10-30 Hz bands use the same low pass as the 1-3 Hz band, since I don't want to see aliasing in the EPICS readouts. I don't think we need faster than 1 Hz readback of the RMS.
  • Confirmed with FOTON that the impulse response for the LP filters are positive for all t >0.

The "C1:PEM-SEIS_STS_1" filter banks are currently empty, so the signal is just in ADC counts. However, by amazing luck, this seems to be the right gain (within a factor of 2) to put the signal into units of microns / second. According the the schematic (D1000749), the default gain of 110 can be switched to make the whole box just have a gain of 2 (differential in, differential out). I wonder if anyone, like Jenne, knows if this is what we have? There's no elog I found about setting the gain switch.

According to the manual, the gain is ~1175 V/(m/s). Our ADC gain should be (2^16)/(40). So:

cal_gain = 1175 * 2 * 65536 / 40  ==>> 0.26 (m/s)/counts

I have put this into the STS_1_X,Y,Z filter modules in c1pem so that these channels are now calibrated. I also put the first few s-domain poles/zeros into the filter based on the manual so that the magnitude in the 10-30 Hz band is correct-ish now.

* Does anyone know how to center the masses on this thing?

Attachment 1: T240_150907.png  83 kB  Uploaded Mon Sep 7 21:02:34 2015  | Show | Show all
ELOG V3.1.3-